OpenStudy (anonymous):

Simplify. square root of 5 over cube root of 5 5 to the power of negative 1 over 6 5 to the power of negative 1 over 6 5 to the power of 5 over 6 im not sure how to do this I thought u were supposed to subtract the exponents but like idk

OpenStudy (anonymous):

\[\huge \frac{ \sqrt{5} }{ \sqrt[3]{5} }?\]

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

\[\sqrt{5} \implies 5^{1/2}\] \[\sqrt[3]{5} \implies 5^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

Yes you're in the right path.

OpenStudy (anonymous):

ok i thought that was how to do it but im not sure how they got the denominator to be 6

OpenStudy (anonymous):

\[\huge \frac{ \sqrt{5} }{ \sqrt[3]{5} } \implies \frac{ 5^{\frac{ 1 }{ 2 }} }{ 5^{\frac{ 1 }{ 3 }} } \implies 5^{\frac{ 1 }{ 2 }-\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }-\frac{ 1 }{ 3 } = ?\]

OpenStudy (anonymous):

yes um -1

OpenStudy (anonymous):

I see the problem, you need to work on your fractions, 1/2 -1/3 is not -1, it's actually 1/6.

OpenStudy (anonymous):

o lol i suck at math but then it would 5 1/6 right

OpenStudy (anonymous):

Correct or \[\sqrt[6]{5}\]

OpenStudy (anonymous):